# UniScene: Unified Occupancy-centric Driving Scene Generation

### v1

@@ 2 Related Work
Semantic Occupancy Representation. Semantic occupancy has emerged as a prominent 3D scene representation in autonomous driving. Current research primarily focuses on Semantic Occupancy Prediction (SOP). MonoScene [11] introduces a 3D SOP method based on a monocular camera. FB-Occ [36] learns features in Bird’s Eye View (BEV) space for SOP. TPVFormer [23] presents a Tri-Perspective View framework for SOP. VPD [31] utilizes a generative diffusion model for 3D SOP. In the realm of occupancy forecasting, OccWorld [86] predicts future occupancy states based on prior occupancy, while OccLlama [68] incorporates Large Language Models (LLMs) to assist in future occupancy prediction. However, research on occupancy generation, particularly for temporal 3D occupancy sequence generation, is still limited. Recently, OccSora [59] employs a Diffusion Transformer (DiT) to generate occupancy, but the quality of the generated results still lags behind the ground truth.
@@ 3.1 Controllable Semantic Occupancy Generation
$\mathcal{L}_{\text{occ}}^{\text{dit}}=\mathbb{E}\left[\sum_{i=1}^{T}\left\|\bm{f}_{\text{dit}}\left(\bm{z}_{\text{occ}}^{i},\mathbf{B}^{i}\right)-\mathcal{N}_{\text{occ}}^{i}(\mathbf{0},\bm{I})\right\|^{2}\right],$ (2)
where $\bm{f}_{\text{dit}}\left(\bm{z}_{\text{occ}}^{i},\mathbf{B}^{i}\right)$ represents the model output, while $\bm{z}_{\text{occ}}^{i}$ and $\mathcal{N}_{\text{occ}}^{i}(\mathbf{0},\bm{I})$ represent the input noisy latent and the target noise of $i^{th}$ frame, respectively. By incorporating temporal information, our occupancy diffusion model enables the generation of long-term consistent occupancy sequences (see Fig. 4 (a)). More architectural details about the Occupancy VAE and DiT can be found in the supplementary.
@@ 3.2 Video: Occupancy as Conditional Guidance
Specifically, the training noise in vanilla noise prior is:
$\bm{\epsilon}_{\text{vid}}^{i}=\lambda\bm{z}_{c}^{i}+\mathcal{N}_{\text{vid}}^{i}(\mathbf{0},\bm{I}),$ (5)
where $\bm{\epsilon}_{\text{vid}}^{i}$ represents the noise input of $i^{th}$ video frame. $\bm{z}_{c}^{i}$ denotes the latent feature of the conditional image frame. $\mathcal{N}_{\text{vid}}^{i}(\mathbf{0},\bm{I})$ and $\lambda$ are the random noise of $i^{th}$ video frame and the balancing coefficient, respectively.
Nevertheless, in real-world scenarios, many regions within dynamic videos exhibit significant variations across multiple frames. The simple strategy described earlier does not account for correspondence modeling in these highly dynamic regions. To address this, we leverage the rendered depth map $\mathbf{D}$ to warp the appearance prior from the reference image to other images using homography warping [10, 67], enabling explicit geometric-aware reprojection. With this approach, we revisit Eq. 5 and optimize it as follows:
$\bm{\epsilon}_{\text{vid}}^{i}=\lambda(\texttt{Warp}(\bm{z}_{c}^{i},\mathbf{D}^{i},\mathbf{K}^{i},\left[\mathbf{R}_{0,i}\mid\mathbf{t}_{0,i}\right]))+\mathcal{N}_{\text{vid}}^{i}(\mathbf{0},\bm{I}),$ (6)
where $\mathbf{K}^{i}$ represent the camera intrinsics and $\left[\mathbf{R}_{0,i}\mid\mathbf{t}_{0,i}\right]$ denote the transform matrix. $\mathbf{D}^{i}$ is the rendered depth map of $i^{th}$ video frame. More details of the evaluation results can be found in Tab. 9 and the supplementary.
@@ Appendix H More Visualization Results
Versatile Generation Ability of UniScene. To further demonstrate the versatile generation capabilities of our proposed UniScene, we apply it across various scenarios. Figure 16 illustrates UniScene’s capacity to generate large-scale, coherent scenes, underscoring the model’s ability to produce outputs that are temporally and spatially consistent in extensive environments. The controllable generation of attribute-diverse videos is showcased in Figure 17, highlighting realistic video outputs with desired attribute variations for different input configurations. Additionally, the application of UniScene to scene editing is depicted in Figure 18, demonstrating the model’s flexibility in generating consistent and realistic outcomes based on edited scene geometries. The results presented in Figures 1920, and 21 further exemplify UniScene’s capability to jointly produce high-quality semantic occupancy, video, and LiDAR data, all while maintaining temporal consistency.

### v2

@@ 2 Related Work
Semantic Occupancy Representation. Semantic occupancy has emerged as a prominent 3D scene representation in autonomous driving. Current research primarily focuses on Semantic Occupancy Prediction (SOP). MonoScene [11] introduces a 3D SOP method based on a monocular camera. TPVFormer [23] presents a Tri-Perspective View framework for SOP. VPD [31] utilizes a generative diffusion model for 3D SOP. In the realm of occupancy forecasting, OccWorld [86] predicts future occupancy states based on prior occupancy, while OccLlama [68] incorporates Large Language Models (LLMs) to assist in future occupancy prediction. However, research on occupancy generation, particularly for temporal 3D occupancy sequence generation, is still limited. Recently, OccSora [59] employs a Diffusion Transformer (DiT) to generate occupancy, but the quality of the generated results still lags behind the ground truth.
@@ 3.1 Controllable Semantic Occupancy Generation
$\mathcal{L}_{\text{occ}}^{\text{dit}}=\mathbb{E}\left[\sum_{i=1}^{T}\left\|\bm{f}_{\text{dit}}\left(\bm{z}_{\text{occ}}^{i},\mathbf{B}^{i}\right)-\bm{\epsilon}_{\text{n}}^{i}\right\|^{2}\right],$ (2)
where $\bm{f}_{\text{dit}}\left(\bm{z}_{\text{occ}}^{i},\mathbf{B}^{i}\right)$ represents the model output, while $\bm{z}_{\text{occ}}^{i}$ represents the input noisy latent of $i^{th}$ frame. $\bm{\epsilon}_{\text{n}}^{i}$ is the random noise sampled from $\mathcal{N}(\mathbf{0},\bm{I})$. By incorporating temporal information, our occupancy diffusion model enables the generation of long-term consistent occupancy sequences (see Fig. 4 (a)). More architectural details about the Occupancy VAE and DiT can be found in the supplementary.
@@ 3.2 Video: Occupancy as Conditional Guidance
$\bm{\epsilon}_{\text{vid}}^{i}=\lambda\bm{z}_{c}+\bm{\epsilon}_{\text{n}}^{i},$ (5)
where $\bm{\epsilon}_{\text{vid}}^{i}$ represents the noise input of $i^{th}$ video frame. $\bm{z}_{c}$ denotes the latent feature corresponding to the conditional video frame, obtained by encoding the conditional video frame using the 3D video VAE encoder. $\bm{\epsilon}_{\text{n}}^{i}$ is random noise sampled from $\mathcal{N}(\mathbf{0},\bm{I})$. $\lambda$ is the balancing coefficient.
Nevertheless, in real-world scenarios, many regions within dynamic videos exhibit significant variations across multiple frames. The simple strategy described earlier does not account for correspondence modeling in these highly dynamic regions. To address this, we leverage the rendered depth map $\mathbf{D}$ to warp the appearance prior from the reference image to other images using depth-based reprojection [10, 67], enabling explicit geometric awareness. With this approach, we revisit Eq. 5 and optimize it as follows:
$\bm{\epsilon}_{\text{vid}}^{i}=\lambda(\texttt{Warp}(\bm{z}_{c},\mathbf{D}^{i},\mathbf{K},\left[\mathbf{R}_{0,i}\mid\mathbf{t}_{0,i}\right]))+\bm{\epsilon}_{\text{n}}^{i},$ (6)
where the depth-based Warp process [10, 67] projects the pixel $p_{i}={(u,v,1)}^{T}$ in the $i^{th}$ latent feature $z_{i}$ to its counterpart ${p}_{0}$ in the conditional latent feature $z_{c}$ as:
${p}_{0}=\mathbf{K}\cdot(\mathbf{R}_{0,i}\cdot({\mathbf{K}}^{-1}\cdot p_{i}\cdot\mathbf{D}^{i}(u,v))+\mathbf{t}_{0,i}).\vspace{-5pt}$ (7)
where $\mathbf{K}$ is the camera intrinsics of the latent features, scaled from the original sensor parameters. $\left[\mathbf{R}_{0,i}\mid\mathbf{t}_{0,i}\right]$ is the transform matrix from target latent $z_{i}$ to conditional latent $z_{c}$. $\mathbf{D}^{i}$ is the rendered depth map of $i^{th}$ video frame. More details of the evaluation results can be found in Tab. 9 and the supplementary.
@@ Appendix F Model Setup
Figure 18: Generalization on the Waymo (upper) and nuPlan (lower) datasets. Due to unavailability, the back view on the Waymo dataset is set to null.
@@ Appendix H More Visualization Results
Generalization on other Autonomous Driving Datasets. As shown in Fig. 18, we evaluate the generalization ability of our model on the Waymo and nuPlan datastes. Given scene layouts, our model can be directly transferred to different datasets with conditional reference images. Note that due to unavailability, the back view on the Waymo dataset is set to null.
@@
Versatile Generation Ability of UniScene. To further demonstrate the versatile generation capabilities of our proposed UniScene, we apply it across various scenarios. Figure 16 illustrates UniScene’s capacity to generate large-scale, coherent scenes, underscoring the model’s ability to produce outputs that are temporally and spatially consistent in extensive environments. The controllable generation of attribute-diverse videos is showcased in Figure 17, highlighting realistic video outputs with desired attribute variations for different input configurations. Additionally, the application of UniScene to scene editing is depicted in Figure 19, demonstrating the model’s flexibility in generating consistent and realistic outcomes based on edited scene geometries. The results presented in Figures 2021, and 22 further exemplify UniScene’s capability to jointly produce high-quality semantic occupancy, video, and LiDAR data, all while maintaining temporal consistency.